# RL-MILP Solver: A Reinforcement Learning Approach for Solving Mixed-Integer Linear Programs with Graph Neural Networks

### v3

@@ 1 Introduction
Existing ML-based primal heuristics have demonstrated their ability to quickly find good solutions (Gasse et al., 2019; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024). However, inaccurate ML predictions hinder solution feasibility, which emphasizes the importance of feasibility (Han et al., 2023). The infeasibility obstructs reaching the optimal solution, posing a significant obstacle to solving MILP. Moreover, the existing studies mostly focus on the prediction for binary decision variables. However, many real-world problems involve non-binary integer (integer, for short) variables such as logistics (Kweon et al., 2024), maritime transportation (Papageorgiou et al., 2014), and energy systems (Ren & Gao, 2010). This highlights the need for techniques that can handle integer variables.
@@
  • We design a RL system that enables the agent to learn problem-solving capabilities by learning relationships between decision variables and constraints.
@@ 2.2.1 Bipartite Graph Representation of MILP
Studies on ML-based primal heuristics represent MILP instances as bipartite graphs (Gasse et al., 2019; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024). In a bipartite graph representation of MILP, one set of nodes is for constraints, and the other is for decision variables. An edge connects a variable node to a constraint node only if the variable appears in the corresponding constraint. For example, in Figure 2(a), the variable $x_{3}$ appears in the constraint $\mathbf{a_{2}}$. Therefore, in the MILP bipartite graph, the variable node for $x_{3}$ is connected to the constraint node for $\mathbf{a_{2}}$.
@@ 3.1.3 Reward
where $\mathcal{F}$ is the feasible region of a given MILP instance, $\alpha=1/\beta$ and $\beta\in(0,1]$ are toward-optimal bias parameters, and $\delta_{t,r}$ and $\delta_{t,b}$ are scaled gaps from $obj_{t+1}$ to the lower bound $obj_{r}$ and the current best objective value $obj_{b}$, respectively.
In $phase1$, the agent aims to find the first feasible solution. Achieving a better (i.e., lower) objective value in $phase1$ sets the agent up for a better starting point in $phase2$. Thus, higher rewards are given as the solution nears $obj_{r}$ (Eq. 12). If the agent finds an infeasible solution $\mathbf{x}_{t+1}\notin\mathcal{F}$, it is corrected by the feasibility reward $\mathcal{R}_{t,\text{F}}$. In $phase2$, the goal is to find feasible solutions better than the current best solution $\mathbf{x}_{b}$. If the agent finds better $\mathbf{x}_{t+1}\in\mathcal{F}$, it earns higher rewards as $\mathbf{x}_{t+1}$ gets closer to $obj_{r}$ (Case 1 in Eq. 13). For $\mathbf{x}_{t+1}\in\mathcal{F}$ worse than $\mathbf{x}_{b}$, the agent receives no reward due to no improvement and is also not penalized to avoid invalidating the learning from in $phase1$ (Case 2 in Eq. 13). If $\mathbf{x}_{t+1}\notin\mathcal{F}$, the agent is encouraged to explore near $\mathbf{x}_{b}$ with smaller $\mathcal{R}_{t,\text{F}}$ as it gets closer to $obj_{b}$ (Case 3, 4 in Eq. 13). According to Proposition 1, we regard $obj_{t+1}<obj_{b}$ as desirable; thus, the penalty is reduced by $\beta$ (Case 3 in Eq. 13). Adjusting $\beta$ further promote exploration toward $obj_{r}$.
@@ 3.1.4 Learning Algorithm
To train the RL agent, we adopt the Actor-Critic (AC) algorithm (Mnih et al., 2016), which has demonstrated effectiveness in solving CO problems (Bello et al., 2016; Hubbs et al., 2020). AC combines a policy-based actor and a value-based critic. The actor optimizes parameters $\theta$ to maximize the expected reward by learning a policy $\pi_{\theta}(\mathcal{A}\mid\mathcal{S})$ that maps the observation $\mathcal{S}$ to a probability distribution over actions $\mathcal{A}$. The critic evaluates $\mathcal{S}$ by learning a value function $V_{\theta}(\mathcal{S})$, which estimates expected return for a given $\mathcal{S}$.
@@ 4.2 Main Results
With a toward-optimal bias (i.e., $\beta=\{0.3,0.6\}$), the compared methods tend to find better solutions. The agent chooses the lesser of two evils, as the potential penalty for solutions with $obj_{t+1}<obj_{b}$ is lower than for those with $obj_{t+1}\geq obj_{b}$ (see Figure 7 in Appendix B). Interestingly, the RL-MILP solver with $\beta=1$ on $D_{9\times 18}$ exhibits the worst Gap among the compared methods, as $\beta=1$ implies no toward-optimal bias. Thus, a suitable toward-optimal bias promotes exploration toward $obj_{r}$, enhancing solution quality. This empirical evidence highlights the importance of guiding the agent’s search, supporting Proposition 1.
@@ 5 Related Work
Bengio et al. (2021) have categorized ML techniques for solving CO into three groups. The first group, Learning to Configure Algorithms, leverages ML to optimize the configuration of specific techniques in traditional solvers. For instance, ML improves solving speed by deciding whether to execute specific operations, such as decomposition (Kruber et al., 2017). The second group, Machine Learning Alongside Optimization Algorithms, integrates ML into traditional solvers to aid key decisions during the optimization process (e.g., selecting the next search region (Lodi & Zarpellon, 2017)). The third group, End-to-End Learning, uses ML to directly learn and predict solutions. This group includes ML-based primal heuristics (Gasse et al., 2019; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024), as well as our method. Existing ML-based primal heuristics output only partial solutions and face challenges in handling non-binary integer variables. They also rely on costly supervised learning for labeled data. A previous RL-based study on MILP focuses solely on finding the first feasible solution (Qi et al., 2021) without optimizing further. To our knowledge, this study is the first full end-to-end RL-based solver that improves solution quality for MILP involving integer variables without relying on traditional solvers.
@@ A.3 Proof
Appendix B Visual Explanation of Reward Function in $phase2$
@@
Figure 7 visualizes the agent’s potential rewards in $phase2$. The second quadrant represents rewards for solutions with $obj_{t+1}<obj_{b}$, which increase as they approach the lower bound $obj_{r}$. The third and fourth quadrants illustrate penalties that increase with the distance of infeasible solutions from the current best objective value $obj_{b}$, assuming $\mathcal{R}_{t,\text{F}}$ as a constant for simplicity. When $obj_{t+1}$ is the same as $obj_{b}$, the penalty becomes zero, but the agent incurs a penalty of -10 only if the agent takes no action (i.e., $\mathbf{x}_{t}=\mathbf{x}_{t+1}$ in Eq. 15). As illustrated in Figures 7(a) and 7(b), a lower $\beta$ (higher $\alpha$) results in higher potential rewards and reduced penalties for solutions with $obj_{t+1}<obj_{b}$. We can promote the agent to take actions that lead to solutions with $obj_{t+1}<obj_{b}$ by controlling the toward-optimal bias parameters $\alpha$ and $\beta$.

### v4

@@ 1 Introduction
Existing ML-based primal heuristics have demonstrated their ability to quickly find good solutions (Ding et al., 2020; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024). However, inaccurate ML predictions hinder solution feasibility, which emphasizes the importance of feasibility (Han et al., 2023). The infeasibility obstructs reaching the optimal solution, posing a significant obstacle to solving MILP. Moreover, the existing studies mostly focus on the prediction for binary decision variables. However, many real-world problems involve non-binary integer (integer, for short) variables such as logistics (Kweon et al., 2024), maritime transportation (Papageorgiou et al., 2014), and energy systems (Ren & Gao, 2010). This highlights the need for techniques that can handle integer variables.
@@
  • We design a RL system that enables the agent to learn problem-solving capabilities by capturing relationships between decision variables and constraints.
@@ 2.2.1 Bipartite Graph Representation of MILP
Studies on ML-based primal heuristics represent MILP instances as bipartite graphs (Ding et al., 2020; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024). In a bipartite graph representation of MILP, one set of nodes is for constraints, and the other is for decision variables. An edge connects a variable node to a constraint node only if the variable appears in the corresponding constraint. For example, in Figure 2(a), the variable $x_{3}$ appears in the constraint $\mathbf{a_{2}}$. Therefore, in the MILP bipartite graph, the variable node for $x_{3}$ is connected to the constraint node for $\mathbf{a_{2}}$.
@@ 3.1.3 Reward
where $\mathcal{F}$ is the feasible region, $\alpha=1/\beta$ and $\beta\in(0,1]$ are toward-optimal bias parameters, and $\delta_{t,r}$ and $\delta_{t,b}$ are scaled gaps from $obj_{t+1}$ to the lower bound $obj_{r}$ and the current best objective value $obj_{b}$, respectively.
In $phase1$, the agent aims to find the first feasible solution. Achieving a better (i.e., lower) objective value in $phase1$ sets the agent up for a better starting point in $phase2$. Thus, higher rewards are given as the solution nears $obj_{r}$ (Eq. 12). If the agent finds an infeasible solution $\mathbf{x}_{t+1}\notin\mathcal{F}$, it is corrected by the feasibility reward $\mathcal{R}_{t,\text{F}}$. In $phase2$, the goal is to find feasible solutions better than the current best solution $\mathbf{x}_{b}$. If the agent finds better $\mathbf{x}_{t+1}\in\mathcal{F}$, it earns higher rewards as $\mathbf{x}_{t+1}$ gets closer to $obj_{r}$ (Case 1 in Eq. 13). For $\mathbf{x}_{t+1}\in\mathcal{F}$ worse than $\mathbf{x}_{b}$, the agent receives no reward due to no improvement and is also not penalized to avoid invalidating the learning from in $phase1$ (Case 2 in Eq. 13). If $\mathbf{x}_{t+1}\notin\mathcal{F}$, the agent receives smaller $\mathcal{R}_{t,\text{F}}$ as it gets closer to $obj_{b}$ (Case 3, 4 in Eq. 13). According to Proposition 1, we regard $obj_{t+1}<obj_{b}$ as desirable; thus, the penalty is reduced by $\beta$ (Case 3 in Eq. 13). Adjusting $\beta$ further promotes exploration toward $obj_{r}$. Appendix B provides further visual explanations of the reward functions.
@@ 3.1.4 Learning Algorithm
We adopt the Actor-Critic (AC) algorithm (Mnih et al., 2016), which has demonstrated effectiveness in solving CO problems (Bello et al., 2016; Hubbs et al., 2020; Zong et al., 2022). AC combines a policy-based actor and a value-based critic. The actor optimizes parameters $\theta$ to maximize the expected reward by learning a policy $\pi_{\theta}(\mathcal{A}\mid\mathcal{S})$ that maps the observation $\mathcal{S}$ to a probability distribution over actions $\mathcal{A}$. The critic evaluates $\mathcal{S}$ by learning a value function $V_{\theta}(\mathcal{S})$, which estimates expected return for a given $\mathcal{S}$.
@@ 4.2 Main Results
With a toward-optimal bias (i.e., $\beta=\{0.3,0.6\}$), the compared methods tend to find better solutions. The agent chooses the lesser of two evils, as the potential penalty for solutions with $obj_{t+1}<obj_{b}$ is lower than for those with $obj_{t+1}\geq obj_{b}$ (see Appendix B.3). Interestingly, the RL-MILP solver with $\beta=1$ on $D_{9\times 18}$ exhibits the worst Gap among the compared methods, as $\beta=1$ implies no toward-optimal bias. Thus, a suitable toward-optimal bias promotes exploration toward $obj_{r}$, enhancing solution quality. This empirical evidence highlights the importance of guiding the agent’s search, supporting Proposition 1.
@@ 5 Related Work
Bengio et al. (2021) have categorized ML techniques for solving CO into three groups. The first group, Learning to Configure Algorithms, leverages ML to optimize the configuration of specific techniques in traditional solvers. For instance, ML improves solving speed by deciding whether to execute specific operations, such as decomposition (Kruber et al., 2017). The second group, Machine Learning Alongside Optimization Algorithms, integrates ML into traditional solvers to aid key decisions during the optimization process (e.g., selecting the next search region (Lodi & Zarpellon, 2017)). The third group, End-to-End Learning, uses ML to directly learn and predict solutions. This group includes ML-based primal heuristics (Ding et al., 2020; Nair et al., 2020; Yoon, 2022; Han et al., 2023; Cantürk et al., 2024), as well as our method. Existing ML-based primal heuristics output only partial solutions and face challenges in handling non-binary integer variables. They also rely on costly supervised learning for labeled data. A previous RL-based study on MILP focuses solely on finding the first feasible solution (Qi et al., 2021) without optimizing further. To our knowledge, this study is the first full end-to-end RL-based solver that improves solution quality for MILP involving integer variables without relying on traditional solvers.
@@ A.3 Proof
Appendix B Visual Explanation of Reward Function
B.1 Feasibility Reward
The feasibility reward $\mathcal{R}_{t,\text{F}}$ is calculated in proportion to the degree of feasibility improvement or deterioration. In Figure 6, the agent violates Constraint-1 while satisfying Constraint-2. If the agent moves closer to the feasible region, it is considered an improvement in feasibility. Conversely, if it moves further away and ends up violating Constraint-2 as well, it is regarded as a deterioration. In this way, rewards are assigned based on how much the agent’s actions improve or worsen feasibility. As a result, the agent learns to find feasible solutions to maximize its rewards.
Figure 6: Illustration of feasibility improvements and deteriorations.
B.2 Optimization Reward in $phase2$
@@
B.3 Toward-Optimal Bias
Figure 8 visualizes the agent’s potential rewards in $phase2$. The second quadrant represents rewards for solutions with $obj_{t+1}<obj_{b}$, which increase as they approach the lower bound $obj_{r}$. The third and fourth quadrants illustrate penalties that increase with the distance of infeasible solutions from the current best objective value $obj_{b}$, assuming $\mathcal{R}_{t,\text{F}}$ as a constant for simplicity. When $obj_{t+1}$ is the same as $obj_{b}$, the penalty becomes zero, but the agent incurs a penalty of -10 only if the agent takes no action (i.e., $\mathbf{x}_{t}=\mathbf{x}_{t+1}$ in Eq. 15). As illustrated in Figures 8(a) and 8(b), a lower $\beta$ (higher $\alpha$) results in higher potential rewards and reduced penalties for solutions with $obj_{t+1}<obj_{b}$. We can promote the agent to take actions that lead to solutions with $obj_{t+1}<obj_{b}$ by controlling the toward-optimal bias parameters $\alpha$ and $\beta$.